# Learnability in Online Kernel Selection with Memory Constraint via Data-dependent Regret Analysis

### v1

@@ Abstract
Online kernel selection is a fundamental problem of online kernel methods. In this paper, we study online kernel selection with memory constraint in which the memory of kernel selection and online prediction procedures is limited to a fixed budget. An essential question is what is the intrinsic relationship among online learnability, memory constraint and data complexity? To answer the question, it is necessary to show the trade-offs between regret bound and memory constraint. Previous work gives a worst-case lower bound depending on the data size, and shows learning is impossible within a small memory constraint. In contrast, we present a different result by providing data-dependent upper bounds depending on two data complexities, namely kernel alignment and the cumulative losses of competitive hypothesis. We propose an algorithmic framework giving data-dependent upper bounds for two types of loss functions. For the hinge loss function, our algorithm achieves an expected upper bound depending on kernel alignment. For smooth loss functions, our algorithm achieves a high-probability upper bound depending on the cumulative losses of competitive hypothesis. We also prove a matching lower bound for smooth loss functions. Our results show that if the two data complexities are sub-linear, then learning is possible within a small memory constraint. Our algorithmic framework depends on a new buffer maintaining framework and a reduction from online kernel selection to prediction with expert advice. Finally, we empirically verify the prediction performance of our algorithms on benchmark datasets.
@@ 1 Introduction
Online kernel selection (OKS) aims to dynamically select a kernel function or a reproducing kernel Hilbert space (RKHS) for online kernel learning algorithms. Compared with offline kernel selection, OKS poses more computational challenges for online kernel selection algorithms due to both the kernel selection procedure and prediction needing to be real-time. We can formulate it as a sequential decision problem. Let $\mathcal{K}=\{\kappa_{i}\}^{K}_{i=1}$ contain $K$ base kernel functions and $\mathcal{H}_{i}$ be a dense subset of the RKHS induced by $\kappa_{i}$. At round $t$, an adversary selects an instances ${\bm{x}}_{t}\in\mathbb{R}^{d}$. Then a learner chooses a hypotheses $f_{t}\in\cup^{K}_{i=1}\mathcal{H}_{i}$ and makes a prediction. The adversary gives the true output $y_{t}$, and the learner suffers a loss $\ell(f_{t}({\bm{x}}_{t}),y_{t})$. We use the regret to measure the performance of the learner which defined as follows,
$\forall\kappa_{i}\in\mathcal{K},\quad\forall f\in\mathcal{H}_{i},~{}\mathrm{Reg}(f)=\sum^{T}_{t=1}\ell(f_{t}({\bm{x}}_{t}),y_{t})-\sum^{T}_{t=1}\ell(f({\bm{x}}_{t}),y_{t}).$
An efficient algorithm should guarantee $\mathrm{Reg}(f)=o(T)$. For convex loss functions, many algorithms [1, 2, 3] achieve (or imply)
@@
The algorithms can adapt to the optimal, yet unknown hypothesis space within a small information-theoretical cost.
A major challenge of OKS is the so called curse of dimensionality, i.e., algorithms must store the previous $t-1$ instances at the $t$-th round. The $O(t)$ memory cost is prohibitive for large-scale online learning problems. To address this issue, we limit the memory of algorithms to a budget of $\mathcal{R}$ quanta. For convex loss functions, the worst-case regret is $\Theta(\max\{\sqrt{T},\frac{T}{\sqrt{\mathcal{R}}}\})$ [4] which proves a trade-off between memory constraint and regret. To be specific, achieving a $O(T^{\alpha})$ regret requires $\mathcal{R}=\Omega(T^{2(1-\alpha)})$, $\alpha\in[\frac{1}{2},1)$. A constant or $\Theta(\ln{T})$ memory cost can not ensure sub-linear regret and thus learning is impossible in any $\mathcal{H}_{i}$. We use the regret to define the learnability of a hypothesis space (Definition 2). However, empirical results have shown that a small memory cost is enough to achieve good prediction performance [5, 6, 7]. It seems that learning is possible in the optimal hypothesis space. To fill the gap between learnability and regret, it is necessary to establish data-dependent regret bounds. To this end, we focus on the following two questions:
@@
    What trade-offs exist between memory constraint and certain data-dependent regret bounds, that is, how do data-dependent regret bounds depend on $\mathcal{R}$, $T$, and $K$?
@@
In this paper, we will answer the two questions affirmatively. We first propose an algorithmic framework. Then we apply the algorithmic framework to two types of loss functions and achieve two kinds of data-dependent regret bounds. To maintain the memory constraint, we reduce it to the budget that limits the size of the buffer. Our algorithmic framework will use the buffer to store a subset of the observed examples. The main results are summarized as follows:
@@
    For the hinge loss function, our algorithm achieves an expected kernel alignment regret as follows (see Theorem 1 for detailed result),
    $\forall\kappa_{i}\in\mathcal{K},~{}\forall f\in\mathbb{H}_{i},~{}\mathbb{E}[\mathrm{Reg}(f)]=\tilde{O}\left(\sqrt{L_{T}(f)\ln{K}}+\frac{\sqrt{\mathcal{R}}}{\sqrt{K}}+\frac{\sqrt{K}}{\sqrt{\mathcal{R}}}\mathcal{A}_{T,\kappa_{i}}\right),$ (1)
@@
    For smooth loss functions (Assumption 3), our algorithm achieves a high-probability small-loss bound (see Theorem 3 for detailed result). $\forall\kappa_{i}\in\mathcal{K},\forall f\in\mathbb{H}_{i}$, with probability at least $1-\delta$,
@@
    where $U>0$ is a constant and bounds the norm of all hypotheses in $\mathbb{H}$. The upper bound in (3) is optimal in terms of the dependence on $L_{T}(f)$.
@@
The data-dependent regret bounds in (1) and (3) improve the $O(\sqrt{T\ln{K}}+\|f\|^{2}_{\mathcal{H}_{i}}\max\{\sqrt{T},\frac{T}{\sqrt{\mathcal{R}}}\})$ [4] bound, since $\mathcal{A}_{T,\kappa_{i}}=O(T)$ and $L_{T}(f)=O(T)$. If $\kappa_{i}$ matches well with the data, then we expect $\mathcal{A}_{T,\kappa_{i}}\ll T$ or $L_{T}(f)\ll T$. In the worst case, i.e., $\mathcal{A}_{T,\kappa_{i}}=\Theta(T)$ or $L_{T}(f)=\Theta(T)$, our bounds are same with previous result. We give new trade-offs between memory constraint and regret bound which answers Q1. For online kernel selection, we just aim to adapt to the optimal kernel $\kappa^{\ast}\in\mathcal{K}$. If $\kappa^{\ast}$ matches well with the data, then we expect $\mathcal{A}_{T,\kappa^{\ast}}=o(T)$ and $\min_{f\in\mathcal{H}_{\kappa^{\ast}}}L_{T}(f)=o(T)$. In this case, a $\Theta(\ln{T})$ memory cost is enough to achieve sub-linear regret which answers Q2. Thus learning is possible in $\mathbb{H}_{\kappa^{\ast}}$ within a small memory constraint.
@@
Our algorithmic framework reduces OKS to prediction with expert advice and uses the well-known optimistic mirror descent framework [8, 9]. We also propose a new buffer maintaining framework.
Related Work
@@
Previous work used buffer maintaining technique and random features to develop algorithms for online kernel selection with memory constraint [1, 3, 7, 10]. However, all of existing results can not answer Q1 and Q2 well. The sketch-based online kernel selection algorithm [7] enjoys a $O(B\ln{T})$ regularized regret (the loss function incorporates a regularizer) within a buffer of size $B$. The regret bound becomes $O(T\ln{T})$ in the case of $B=\Theta(T)$. The Raker algorithm [10] achieves a $O(\sqrt{T\ln{K}}+T/\sqrt{D})$  [^1] The original regret bound is $O(U\sqrt{T\ln{K}}+\epsilon TU)$, and holds with probability $1-\Theta\left(\epsilon^{-2}\exp(-D\epsilon^{2}/(4d+8))\right)$. regret and suffers a space complexity in $O((d+K)D)$, where $D$ is the number of random features. The upper bound also shows a trade-off in the worst case. If $D$ is a constant, then Raker can not achieve a sub-linear regret bound which gives a negative answer to Q2.
@@
Our work is also related to achieving data-dependent regret bounds for online learning and especially online kernel learning. For online learning, various data-dependent regret bounds including small-loss bound [11, 12, 13], variance bound [14, 15] and path-length bound [8, 16, 17], have been achieved. The adaptive online learning framework [18] can achieve data-dependent and model-dependent regret bounds, but can not induce a computationally efficient algorithm. For online kernel learning, achieving data-dependent regret bounds is harder, since we must balance the computational cost. For loss functions satisfying specific smoothness (Assumption 3), the OSKL algorithm [19] achieves a small-loss bound. For loss functions with a curvature property, the PROS-N-KONS algorithm [20] achieves a regret bound depending on the effective dimension of the kernel matrix. All of these algorithms did not consider memory constraint, and thus did not give a trade-off between memory constraint and regret bound.
@@ 2 Problem Setup
the Bregman divergence associated with a strongly convex regularizer $\psi:\mathcal{H}_{\kappa}\rightarrow\mathbb{R}$.
If an oracle gives the optimal kernel $\kappa^{\ast}\in\mathcal{K}$ for $\mathcal{I}_{T}$, then we just run an online kernel learning algorithm in $\mathcal{H}_{\kappa^{\ast}}$. Since $\kappa^{\ast}$ is unknown, the learner aims to develop a kernel selection algorithm and generate a sequence of hypotheses $\{f_{t}\}^{T}_{t=1}$ which is competitive to any $f\in\mathcal{H}_{\kappa^{\ast}}$, that is, the learner hopes to guarantee $\mathrm{Reg}(f)=o(T)$. The representer theorem proves that $f=\sum^{T}_{t=1}a_{t}\kappa_{i}({\bm{x}}_{t},\cdot)$ for any $f\in\mathcal{H}_{i}$. Thus storing $f_{t}$ needs to store $\{({\bm{x}}_{\tau},y_{\tau})\}^{t-1}_{\tau=1}$ and leads to a memory cost in $O(t)$. In this paper, we consider online kernel selection with memory constraint. Next we define the memory constraint and reduces the memory constraint to the size of a buffer.
Definition 1 (Memory Budget [4]).
@@
Assumption 1 implies that there is no need to assume $\mathcal{R}=\infty$ unless $T=\infty$, since the maximal $\mathcal{R}$ satisfies $B=T$. We define budgeted online kernel learning algorithms as such algorithms that operate on a subset of the observed examples, such as Forgetron [5], BSGD [21] to name but a few. In Assumption 1, $\alpha$ is independent of kernel function, since the memory cost is used to store the examples and the coefficients.
@@
There is a positive constant $k_{1}$, such that $\forall\kappa_{i}\in\mathcal{K}$ and $\forall{\bm{u}}\in\mathcal{X}$, $\kappa_{i}({\bm{u}},{\bm{u}})\in[k_{1},1]$.
We will define two types of data complexities and use them to bound the regret. The first one is called kernel alignment which is defined as follows
@@
where ${\bm{Y}}=(y_{1},\ldots,y_{T})^{\mathrm{T}}$ and ${\bm{K}}_{i}$ is the kernel matrix induced by $\kappa_{i}$. The measure $\mathcal{A}_{T,\kappa_{i}}$ quantifies the matching between ${\bm{K}}_{i}$ and the examples. If ${\bm{K}}_{i}={\bm{Y}}{\bm{Y}}^{\mathrm{T}}$ is the ideal kernel matrix, then $\mathcal{A}_{T,\kappa_{i}}=\Theta(1)$. The second one is called small-loss defined by $\min_{f\in\mathbb{H}_{i}}L_{T}(f)$ where $L_{T}(f)$ follows (2) and
@@
Finally, we define the online learnability which is a variant of the learnability defined in [22].
@@
The definition is equivalent to $\mathrm{Reg}(f)=o(T)$. We can also replace $\mathrm{Reg}(f)$ with $\mathbb{E}[\mathrm{Reg}(f)]$. The learnability defined in [22] considers the whole sample space, i.e., the worst-case examples, while our definition only considers a fixed $\mathcal{I}_{T}$. In the worst-case, the two kinds of learnability are equivalent. Since $\mathcal{I}_{T}$ may not always be generated in the worst-case, our learnability can adapt to the hardness of $\mathcal{I}_{T}$.

### v2

@@ Abstract
Online kernel selection is a fundamental problem of online kernel methods. In this paper, we study online kernel selection with memory constraint in which the memory of kernel selection and online prediction procedures is limited to a fixed budget. An essential question is what is the intrinsic relationship among online learnability, memory constraint, and data complexity? To answer the question, it is necessary to show the trade-offs between regret and memory constraint. Previous work gives a worst-case lower bound depending on the data size, and shows learning is impossible within a small memory constraint. In contrast, we present distinct results by offering data-dependent upper bounds that rely on two data complexities: kernel alignment and the cumulative losses of competitive hypothesis. We propose an algorithmic framework giving data-dependent upper bounds for two types of loss functions. For the hinge loss function, our algorithm achieves an expected upper bound depending on kernel alignment. For smooth loss functions, our algorithm achieves a high-probability upper bound depending on the cumulative losses of competitive hypothesis. We also prove a matching lower bound for smooth loss functions. Our results show that if the two data complexities are sub-linear, then learning is possible within a small memory constraint. Our algorithmic framework depends on a new buffer maintaining framework and a reduction from online kernel selection to prediction with expert advice. Finally, we empirically verify the prediction performance of our algorithms on benchmark datasets.
@@ 1 Introduction
Online kernel selection (OKS) aims to dynamically select a kernel function or a reproducing kernel Hilbert space (RKHS) for online kernel learning algorithms. Compared to offline kernel selection, OKS poses more computational challenges for kernel selection algorithms, as both the kernel selection procedure and prediction need to occur in real-time. We can formulate it as a sequential decision problem. Let $\mathcal{K}=\{\kappa_{i}\}^{K}_{i=1}$ contain $K$ base kernel functions and $\mathcal{H}_{i}$ be a dense subset of the RKHS induced by $\kappa_{i}$. At round $t$, an adversary selects an instances ${\bm{x}}_{t}\in\mathbb{R}^{d}$. Then a learner chooses a hypothesis $f_{t}\in\cup^{K}_{i=1}\mathcal{H}_{i}$ and makes a prediction. The adversary gives the true output $y_{t}$, and the learner suffers a loss $\ell(f_{t}({\bm{x}}_{t}),y_{t})$. We use the regret to measure the performance of the learner which is defined as
$\forall\kappa_{i}\in\mathcal{K},\quad\forall f\in\mathcal{H}_{i},\quad\mathrm{Reg}(f)=\sum^{T}_{t=1}\ell(f_{t}({\bm{x}}_{t}),y_{t})-\sum^{T}_{t=1}\ell(f({\bm{x}}_{t}),y_{t}).$
An efficient algorithm should guarantee $\mathrm{Reg}(f)=o(T)$. For convex loss functions, many algorithms [Sahoo2014Online, Foster2017Parameter, Liao2021High] achieve (or imply)
@@
The algorithms adapt to the optimal, yet unknown hypothesis space within a small information-theoretical cost.
A major challenge in OKS is the so called curse of dimensionality, that is, algorithms must store the previous $t-1$ instances at the $t$-th round. The $O(t)$ memory cost is prohibitive for large-scale online learning problems. To address this issue, we limit the memory of algorithms to a budget of $\mathcal{R}$ quanta. For convex loss functions, the worst-case regret is $\Theta\left(\max\left\{\sqrt{T},\frac{T}{\sqrt{\mathcal{R}}}\right\}\right)$ [Li2022Worst] which proves a trade-off between memory constraint and regret. To be specific, achieving a $O(T^{\alpha})$ regret requires $\mathcal{R}=\Omega\left(T^{2(1-\alpha)}\right)$, $\alpha\in\left[\frac{1}{2},1\right)$. Neither a constant nor a $\Theta(\ln{T})$ memory cost can guarantee sub-linear regret, thereby learning is impossible in any $\mathcal{H}_{i}$. We use the regret to define the learnability of a hypothesis space (see Definition 2). However, empirical results have shown that a small memory cost is enough to achieve good prediction performance [Dekel2008The, Zhao2012Fast, Zhang2018Online]. It seems that learning is possible in the optimal hypothesis space. To fill the gap between learnability and regret, it is necessary to establish data-dependent regret bounds. To this end, we focus on the following two questions:
@@
    What are the new trade-offs between memory constraints and regret, that is, how do certain data-dependent regret bounds depend on $\mathcal{R}$, $T$, and $K$?
@@
In this paper, we will answer the two questions affirmatively. We first propose an algorithmic framework. Then we apply the algorithmic framework to two types of loss functions and achieve two kinds of data-dependent regret bounds. To maintain the memory constraint, we reduce it to the budget that limits the size of the buffer. Our algorithmic framework will use the buffer to store a subset of the observed examples. The main results are summarized as follows.
@@
    For the hinge loss function, our algorithm enjoys an expected kernel alignment regret bound as follows (see Theorem 1 for detailed result),
    $\forall\kappa_{i}\in\mathcal{K},\forall f\in\mathbb{H}_{i},\quad\mathbb{E}[\mathrm{Reg}(f)]=\tilde{O}\left(\sqrt{L_{T}(f)\ln{K}}+\frac{\sqrt{\mathcal{R}}}{\sqrt{K}}+\frac{\sqrt{K}}{\sqrt{\mathcal{R}}}\mathcal{A}_{T,\kappa_{i}}\right),$ (1)
@@
    For smooth loss functions (see Assumption 3), our algorithm achieves a high-probability small-loss bound (see Theorem 3 for detailed result). $\forall\kappa_{i}\in\mathcal{K},\forall f\in\mathbb{H}_{i}$, with probability at least $1-\delta$,
@@
    in which $U>0$ is a constant and bounds the norm of all hypotheses in $\mathbb{H}$. The upper bound in (3) is optimal in terms of the dependence on $L_{T}(f)$.
@@
The data-dependent bounds in (1) and (3) improve the bound $O\left(\sqrt{T\ln{K}}+\|f\|^{2}_{\mathcal{H}_{i}}\max\{\sqrt{T},\frac{T}{\sqrt{\mathcal{R}}}\}\right)$ [Li2022Worst], given that $\mathcal{A}_{T,\kappa_{i}}=O(T)$ and $L_{T}(f)=O(T)$. If $\kappa_{i}$ matches well with the data, then we expect $\mathcal{A}_{T,\kappa_{i}}\ll T$ or $L_{T}(f)\ll T$. In the worst case, i.e., $\mathcal{A}_{T,\kappa_{i}}=\Theta(T)$ or $L_{T}(f)=\Theta(T)$, our bounds are same with previous result. We give new trade-offs between memory constraint and regret which answers Q1. For online kernel selection, we just aim to adapt to the optimal kernel $\kappa^{\ast}\in\mathcal{K}$. If $\kappa^{\ast}$ matches well with the data, then we expect $\mathcal{A}_{T,\kappa^{\ast}}=o(T)$ and $\min_{f\in\mathcal{H}_{\kappa^{\ast}}}L_{T}(f)=o(T)$. In this case, a $\Theta(\ln{T})$ memory cost is enough to achieve sub-linear regret which answers Q2. Thus learning is possible in $\mathbb{H}_{\kappa^{\ast}}$ within a small memory constraint.
@@
Our algorithmic framework reduces OKS to prediction with expert advice and uses the well-known optimistic mirror descent framework [Chiang2012Online, Rakhlin2013Online]. We also propose a new buffer maintaining framework.
2 Related Work
@@
Previous work has adopted buffer maintaining technique and random features to develop algorithms for online kernel selection with memory constraint [Sahoo2014Online, Liao2021High, Zhang2018Online, Shen2019Random]. However, all of existing results can not answer Q1 and Q2 well. The sketch-based online kernel selection algorithm [Zhang2018Online] enjoys a $O(B\ln{T})$ regularized regret (the loss function incorporates a regularizer) within a buffer of size $B$. The regret bound becomes $O(T\ln{T})$ in the case of $B=\Theta(T)$. The Raker algorithm [Shen2019Random] achieves a $O\left(\sqrt{T\ln{K}}+\frac{T}{\sqrt{D}}\right)$  [^1] The original regret bound is $O(U\sqrt{T\ln{K}}+\epsilon TU)$, and holds with probability $1-\Theta\left(\epsilon^{-2}\exp\left(-\frac{D}{4d+8}\epsilon^{2}\right)\right)$. regret and suffers a space complexity in $O((d+K)D)$, where $D$ is the number of random features. The upper bound also shows a trade-off in the worst case. If $D$ is a constant or $D=\Theta(\ln{T})$, then Raker can not achieve a sub-linear regret bound, providing a negative answer to Q2.
@@
Our work is also related to achieving data-dependent regret bounds for online learning and especially online kernel learning. For online learning, various data-dependent regret bounds including small-loss bound [Cesa-Bianchi2006Prediction, Lykouris2018Small, Lee2020Bias], variance bound [Hazan2009Better, Hazan2010Extracting] and path-length bound [Chiang2012Online, Steinhardt2014Adaptivity, Wei2018More], have been established. The adaptive online learning framework [Foster2015Adaptive] can achieve data-dependent and model-dependent regret bounds, but can not induce a computationally efficient algorithm. For online kernel learning, it is harder to achieve data-dependent regret bounds, as we must balance the computational cost. For loss functions satisfying specific smoothness (see Assumption 3), the OSKL algorithm [Zhang2013Online] achieves a small-loss bound. For loss functions with a curvature property, the PROS-N-KONS algorithm [Calandriello2017Efficient] and the PKAWV algorithm [Jezequel2019Efficient] achieve regret bounds depending on the effective dimension of the kernel matrix. None of these algorithms took memory constraints into account, failing to provide a trade-off between memory constraint and regret.
@@ 3 Problem Setup
the Bregman divergence associated with a strongly convex regularizer $\psi(\cdot):\mathcal{H}_{\kappa}\rightarrow\mathbb{R}$.
Let $\mathcal{K}=\{\kappa_{i}\}^{K}_{i=1}$ contain $K$ base kernels. If an oracle gives the optimal kernel $\kappa^{\ast}\in\mathcal{K}$ for $\mathcal{I}_{T}$, then we just run an online kernel learning algorithm in $\mathcal{H}_{\kappa^{\ast}}$. Since $\kappa^{\ast}$ is unknown, the learner aims to develop a kernel selection algorithm and generate a sequence of hypotheses $\{f_{t}\}^{T}_{t=1}$ that can compete with any $f\in\mathcal{H}_{\kappa^{\ast}}$, that is, the learner hopes to guarantee $\mathrm{Reg}(f)=o(T)$. For any $f\in\mathcal{H}_{i}$, we have $f=\sum^{T}_{t=1}a_{t}\kappa_{i}({\bm{x}}_{t},\cdot)$. Thus storing $f_{t}$ needs to store $\{({\bm{x}}_{\tau},y_{\tau})\}^{t-1}_{\tau=1}$, incurring a memory cost in $O(t)$. In this paper, we consider online kernel selection with memory constraint. Next we define the memory constraint and reduce it to the size of a buffer.
Definition 1 (Memory Budget [Li2022Worst]).
@@
Assumption 1 implies that there is no need to assume $\mathcal{R}=\infty$ unless $T=\infty$, as the maximal value of $\mathcal{R}$ satisfies $B=T$. The budgeted online kernel learning algorithms are such algorithms that operate on a subset of the observed examples, such as Forgetron [Dekel2008The], BSGD [Wang2012Breakingecond] to name but a few. In Assumption 1, $\alpha$ is independent of kernel function, since the memory cost is used to store the examples and the coefficients.
@@
There is a constant $k_{1}>0$, such that $\forall\kappa_{i}\in\mathcal{K}$ and $\forall{\bm{u}}\in\mathcal{X}$, $\kappa_{i}({\bm{u}},{\bm{u}})\in[k_{1},1]$.
We will define two types of data complexities and use them to bound the regret. The first one is called kernel alignment defined as follows
@@
where ${\bm{Y}}=(y_{1},\ldots,y_{T})^{\mathrm{T}}$ and ${\bm{K}}_{i}$ is the kernel matrix induced by $\kappa_{i}$. $\mathcal{A}_{T,\kappa_{i}}$ quantifies how well ${\bm{K}}_{i}$ matches the examples. If ${\bm{K}}_{i}={\bm{Y}}{\bm{Y}}^{\mathrm{T}}$ is the ideal kernel matrix, then $\mathcal{A}_{T,\kappa_{i}}=\Theta(1)$. The second one is called small-loss defined by $\min_{f\in\mathbb{H}_{i}}L_{T}(f)$ where $L_{T}(f)$ follows (2) and
@@
Finally, we define the online learnability which is a variant of the learnability defined in [Rakhlin2010Online].
@@
The definition is equivalent to $\mathrm{Reg}(f)=o(T)$. We can also replace $\mathrm{Reg}(f)$ with $\mathbb{E}[\mathrm{Reg}(f)]$. The learnability defined in [Rakhlin2010Online] holds for the entire sample space, i.e., the worst-case examples, while our definition only considers a fixed $\mathcal{I}_{T}$. In the worst-case, the two kinds of learnability are equivalent. Given that $\mathcal{I}_{T}$ may not always be generated in the worst-case, our learnability can adapt to the hardness of $\mathcal{I}_{T}$.